show:
cot^2x/1-cos^2x=csc^2x

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

show:
cot^2x/1-cos^2x=csc^2x
--------
If you mean:
cot^2(x)/(1-cos^2(x))=csc^2(x)
cot^2(x)/(sin^2(x))=csc^2(x)
Multiply by sin^2(x)
cot^2(x) = 1
------
That can be solved for x, but you said "show."


Answer by solver91311(24713)



It looks like you mean:



And that you are asking to show that this is an identity.

This is unfortunate, because the equation shown is not an identity.

It is only true when which is only true when

















John

My calculator said it, I believe it, that settles it


I > Ø